A 2.20g sample of K2Cr2O7 is dissolved in about 50mL of water and the solution made up to 500.0mL in a volumetric flask.
(i). Calculate the concentration of the K2Cr2O7 solution.
(ii). A 20.0mL aliquot of this solution is diluted to 100.0mL. What is the concentration of this solution?
(iii). What volume of the initial solution would be required to make up 100.0mL of
7.00 X 10-3 mol L-1 K2Cr2O7 solution?
1
Expert's answer
2020-08-02T15:52:12-0400

i) Equation for the reaction is;

"K_2Cr_2O_7+H_2O\to2KCrO_4+2H"

Molar mass of "K_2Cr_2O_7=294.1846"

Molar mass of "H_2O= 18. 01528"

Concentration of "K_2Cr_2O_7" solution

Molarity"=" "Moles of solute\over Liter of solution"

"= 2.20gK_2Cr_2O_7\over500mL" "*" "1mol K_2Cr_2O_7 \over 294. 1846gK_2Cr_2O_7L" "*" "50nL\over L"

"=" "3.748*10^{-3}" "mol\/L"

ii) "2.20gK_2Cr_2O_7\over 100mL" "*" "2mol \over 294.1846K_2Cr_2O_7" "*" "20nO\over L"

"= 2.99*10^{-3}mol\/L"

iii) "Volume=" "Mass\over Concentration"

"\therefore 7.00*10^{-3}" "=" "x\over 2.99*10^{-3}"

"x=2.093*10^{-5}"

Vol of solution"=" "2.093*10^{-5}L"